In a right triangle the sum of the 2 legs is 23. The hypotenuse is 2 more than the longer leg. What are the dimensions of the triangle?
Answer by mananth(16946)


(x+y) = 23
x^2+y^2= (y+2)^2
x^2+y^2=y^2+4y+4
x^2=4y+4
but y=23-x
x^2=4(23-x)+4
x^2=92-4x+4
x^2+4x-96=0
Solve quadratic equation to get x
x=-12 OR 8
Ignore negative
x= 8
y = 23-8= 15
hypotenuse = 17